# Few-Shot Retinomorphic Vision in a Nonlinear Photonic Network Laser

### v4

@@ Introduction
Single-layer classification (fig. 3a) works by projecting whole images onto the network, measuring the hyperspectrum, then applying logistic regression on the training set hyperspectra to learn a fixed weight (per image class) for each channel of the hyperspectral output. Once the system is trained, unseen images are classified by performing a weighted sum over the hyperspectral channel amplitudes and learned weights for each class. The class with the largest weighted sum is returned as the network prediction. The single-layer scheme obtains accuracy of 96.03% on MNIST digits, 85.18% on Fashion MNIST, and 86.59% on BreaKHis cancer diagnosis, with measurement throughput up to 100 Hz.
@@
Multi-layer classification begins with feature detection (fig. 3b). Ten spectral regions performing feature detection were selected, producing ten feature maps per input image. Regions returning clear edge-detection were hand selected, with future improvement available by training the feature selection process. Three feature maps are shown schematically in figure 3b), with all feature maps for a range of MNIST digits and Fashion-MNIST shown in supplementary information. These feature maps are then re-illuminated on the InP network as whole images and their corresponding hyperspectra recorded, illustrated in figure 3c). The hyperspectral output from the second-layer feature map illuminations are combined in parallel and a single logistic regression step is applied to produce classification results. The multi-layer scheme obtains accuracy of 98.05% on MNIST digits, 87.85% on Fashion-MNIST, and 90.12% on BreaKHis cancer diagnosis. Further details of the regression process are given in the Methods.
b) Pie-chart illustrating the heavy class-imablance in the dataset.
@@
c) Class-balanced accuracy vs. training set length, comparing photonic network against software benchmarks. Test set is fixed at 1015 images. Class-balanced accuracy penalises incorrect classification on rarer classes more heavily, ensuring performance is not dominated by majority classes. Photonic network scores 53.7% max accuracy, beyond software models.
@@
Network laser performance relative to software benchmarks is assessed in figures 3g,h,i). Classification accuracy vs. training set length is assessed for single- and multi-layer network laser schemes against a range of software neural networks. Five software approaches are considered: logistic regression on the raw images (a direct control, effectively bypassing our physical system), a multi-layer perceptron (300100 neurons), a LeNet5 CNN (22 filters), a larger CNN (96 filters), and the much more powerful ‘EfficientNetV2-B0’[40, 41] CNN (3040 filters, 7.9 million parameters) using a transfer learning approach - initialised with pre-trained ImageNet weights, which are then updated during training. Full details of the software networks are provided in the Methods. We also benchmark against a ‘quasi-isomorphic’ software network: an initial convolutional filter layer of 10 44 filters followed by 30,000 randomly weighted ReLu neurons and a final logistic regression layer, mirroring the structure of our system. Results are shown in supplementary information.
Few-shot/low-data performance The photonic network laser excels when training data is scarce, a regime where conventional AI can struggle. The network outperforms all considered software models on training sets smaller than 5,000 images (digits), 15,000 images (fashion), and across all training lengths on the harder biomedical BreaKHis task — including the powerful EfficientNet CNN. We score 77.7% accuracy on BreaKHis from just 10 training images (5 examples per class), far higher than the software models. This highlights the effective learning provided by the nonlinear photonic dynamics and mode competition of the network laser. Here we compare performance when using only the raw training set. Software approaches can achieve higher scores on the BreaKHis task, but this requires costly extra steps including artificial augmentation of the training set, and extensive fine-tuning of model weights.
@@
The neuromorphic capabilities of the photonic network are not restricted to feature detection and classification. Here, we demonstrate the ability of the network to spatially map regions of interest in biomedical images - drawing masks over skin lesions in colour photographic images from the HAM10000/ISIC 2018 dataset[44, 43] (fig. 4a, 10015 images of seven different classes of skin lesion). This task, termed segmentation, demonstrates that while the hyperspectral output of our network only contains information on the wavelength and one-dimensional spatial position of the lasing response, it is capable of physically encoding and processing complex two-dimensional spatial information - required for segmentation and other spatial processing. The task here is challenging: lesions are non-uniform in shape, a variety of colours and intensities, and many images contain darker non-lesion regions such as hairs, moles, and areas of uneven lighting. We combine two distinct tasks on the HAM10k dataset: class-imbalanced diagnostic classification, separating lesions into seven classes, and segmentation, drawing spatial masks to indicate lesion location. This two-fold processing has potential future impact in automated/semi-automated combined diagnosis and surgery systems. We need only input the HAM10k images to our photonic network once to achieve both tasks, simply retraining our computationally-cheap linear output regression layer to change between diagnosis and segmentation.
@@
Class-imbalanced diagnosis Here we diagnose lesion images into seven classes. This is made highly challenging by the heavy class-imbalance of the dataset - illustrated in figure 4b). Machine learning typically struggles with class imbalance, tending to incorrectly classify rare classes as examples of the more common classes. Overcoming this requires costly techniques such as data augmentation, e.g. generating many distorted copies of training examples and using them to artificially extend the training set. The HAM10000 task is heavily imbalanced, with the ‘NV’ class representing 66.9% of the dataset, and four of the seven classes each representing 5.1% or less. Figure 4c) shows class-balanced classification accuracy against training set length (class-balanced accuracy penalises incorrect guesses on rare classes more heavily), with the photonic network scoring 53.7% max accuracy, outperforming all software benchmarks at all training lengths. This shows our system demonstrates strong advantages on both few-shot learning and heavily-imbalanced datasets - a hard problem.
@@
Biomedical segmentation For segmentation, we use the same hyperspectral responses from the lesion images which we use for classification. Here, rather than using the diagnostic class as training/testing labels, and logistic regression as an output layer - we use the ground-truth segmentation maps as training/testing labels and ridge regression as an output layer (as the segmentation task has continuous output, rather than discrete labels). Otherwise the procedure follows that used for classification. Figure 4d) compares three segmentation metric scores across the test set: DICE (F1 score for segmentation), Jaccard (intersection over union) and pixel accuracy (correct pixels over incorrect). DICE and Jaccard provide best indication of segmentation quality, pixel accuracy scores tend to be distorted by whether the correct ground-truth mask occupies a small or large fraction of the total image. The photonic network scores highest on all metrics, highlighting the ability of the network to implement spatial processing.
@@ Conclusions
This work introduces a new class of photonic hardware for compact, highly nonlinear neuromorphic processing which excels at few-shot learning — with potential impact on biomedical sensing, edge computing, and beyond. The system performs parallel feature detection, heterogenous nonlinear activation, and strong image classification and segmentation in a single 150 m on-chip InP network, outperforming powerful modern software CNNs such as EfficientNet particularly on short and imbalanced datasets.
We provide routes to solutions for both the physical nonlinearity and large footprint issues facing many existing photonic computing schemes. As the lasing modes responsible for feature detection spatially overlap, we provide compact feature detection of multiple image features in a micron-scale footprint. In contrast, existing physical convolutional hardware schemes typically detect one feature per physical repeat of the device, hence the spatial footprint, fabrication complexity and energy consumption typically scale with the number of features detected.
Strong learning in the few-shot/low-data regime with scarce training examples is a hard and open challenge in machine learning. The performance of our photonic network in this regime, including on hard biomedical tasks such as BreaKHis and HAM10k, provides valuable evidence that physics-based neuromorphic systems can handle complex, data-constrained learning tasks. These results bring physical neuromorphic schemes closer to offering viable alternatives to conventional software-based approaches. Our good performance here extends to heavily class-imbalanced datasets, which are similarly challenging for conventional software AI. One of the key proposed applications for neuromorphic computing hardware is ‘edge computing’ where remotely deployed hardware performs computation directly at the point of sensing, adapting on-the-fly to incoming data and changing environmental conditions without access to large offline/cloud-based training data. Here, it is highly advantageous to perform well with short and imbalanced training sets, as locally collected data will typically be far from the ideal large and class-balanced training sets where software AI is most comfortable.
@@
RS, KM, HS, J.D., AF, DS, and W.K.N. designed and developed the network laser concept.
@@
JCG wrote the manuscript with contributions from all authors.
@@ Acknowledgements
This work was supported by the Royal Academy of Engineering Research Fellowships, awarded to JCG.
@@
TVR, DS, and RS acknowledge support from The Engineering and Physical Sciences Research Council (EPSRC), grant number EP/T027258.
WKN acknowledges the research support funded by the President’s PhD Scholarships from Imperial College London.
This work was supported by the EPSRC Impact Acceleration Account fund awarded to JCG and RS.
This work was supported by the Imperial College London DT-Prime scheme awared to JCG, RS and MB.
@@
TF was supported by the EPSRC doctoral training programme PhD award.
@@
We thank David Mack for excellent laboratory management and Stephen Cussell and David Bowler for excellent technical workshop services at Imperial College London.
@@
Test image of ‘Mandrill’ was taken from the USC-SIPI image database, available at:
https://sipi.usc.edu/database/database.php
@@
The authors declare no competing interests.
Data availability statement
The datasets generated during and/or analysed during the current study are available from the corresponding author on reasonable request.
Code availability statement
The code used in this study is available from the corresponding author on reasonable request.
@@ Methods
The network was designed by using a randomly-distributed point grid pattern on a plane and generating a Voronoi diagram[32]. The network is cropped to a circular area with 150 m diameter, and the density of the point patterns is varied to generate waveguide graph-edges with 350 nm width and an average length of 5 m.
@@
The edge-detection and digit classification experiments are performed in an optical microscopy setup with a spatially-structured pump (via DMD) to excite the network laser structure. A schematic of the setup is shown in the supplementary information). The pump laser is a 633 nm femtosecond pulsed laser (Light Conversion PHAROS with ORPHEUS-VIS Optical Parametric Amplifier (OPA) and LYRA harmonic generator, 200 fs pulse duration, 1 kHz repetition rate) is first expanded through a telescope and shaped by reflecting on a digital micromirror device (DMD, ViALUX V-650L). The DMD consist of programmable mirror pixels which shape the incident uniform beam into any pixelated black-and-white image (i.e. training images). The power of the shaped pump patterns is modulated by a motorised continuously variable neutral density (ND) filter, and the exact power received on the sample is monitored by a power meter (Thorlabs, S120VC Si photodiode, 200 - 1100 nm) on the split path reflected from a glass slide inserted. The spatially-structured pump is focused on the sample through a objective (Olympus Plan N 20x with a 0.4 numerical aperture (NA)). The lasing emission from the network is collected through the same objective, filtered (Thorlabs FELH0750, long-pass filter), and spatially and spectrally analysed using a grating spectrometer (Princeton Instruments SpectraPro HRS-300) equipped with a g/mm visible grating without an entrance slit. The signals are measured by a 2D charge-coupled device camera (CCD, Princeton Instruments Pixis 256) with a spatial resolution of m.
@@
The lasing modes and intensities of the network are modelled using netSALT[26]. The designed network is represented as a graph, with the edges subdivided to overlap with an 88 grid to correspond to the kernel. When the network is unpumped, each edge has a complex refractive index n = 3.4 + 0.001i. The edges are further assumed to have a Lorentzian gain profile centered at normalised frequency with a width , and are pumped up to a normalised power (For a detailed physical explanation of the parameters, see Ref. 26). A separate simulation is performed for each 88 binary kernel, by pumping only the edges falling under the illuminated pixels. The passive (without pumping) modes of the network are first identified in the complex frequency plane , and these are the same for all kernels. For each kernel, the modes are tracked in the complex plane as the pump power is increased, allowing to compare between the same modes across different kernels. As a mode reaches the real frequency plane [Im], it attains threshold and starts lasing. Beyond this pump power, the frequency and shape of the mode do not change but mode competition between the lasing modes modifies their lasing intensities. They further deplete the gain available to the remaining modes, increasing their thresholds. The mode frequencies, thresholds and intensities found by netSALT incorporate the mode competition, gain depletion and nonlinearity.

### v5

@@ Introduction
Single-layer classification (fig. 3a) works by projecting whole images onto the network, measuring the hyperspectrum, then applying logistic regression on the training set hyperspectra to learn a fixed weight (per image class) for each channel of the hyperspectral output. This feed-forward single layer is similar in some respects to an Extreme Learning Machine, where input information is nonlinearly transformed and projected to a high dimensional state by a fixed set of complex, nonlinear nodes, and then transformed to a useful computational output via a linear readout layer[54]. However, the ability of the network laser to directly produce computationally relevant outputs such as image feature maps from the intrinsic nonlinear photonic dynamics, without requiring a trained linear readout layer, means that our system goes beyond the definition of an extreme learning machine. Once the system is trained, unseen images are classified by performing a weighted sum over the hyperspectral channel amplitudes and learned weights for each class. The class with the largest weighted sum is returned as the network prediction. The single-layer scheme obtains accuracy of 96.03% on MNIST digits, 85.18% on Fashion MNIST, and 86.59% on BreaKHis cancer diagnosis, with measurement throughput up to 100 Hz.
@@
Multi-layer classification begins with feature detection (fig. 3b). Ten spectral regions performing feature detection were selected, producing ten feature maps per input image. Regions returning clear edge-detection were hand selected, with future improvement available by training the feature selection process. Three feature maps are shown schematically in figure 3b), with all feature maps for a range of MNIST digits and Fashion-MNIST shown in supplementary figures S15 and S16. These feature maps are then re-illuminated on the InP network as whole images and their corresponding hyperspectra recorded, illustrated in figure 3c). The hyperspectral output from the second-layer feature map illuminations are combined in parallel and a single logistic regression step is applied to produce classification results. The multi-layer scheme obtains accuracy of 98.05% on MNIST digits, 87.85% on Fashion-MNIST, and 90.12% on BreaKHis cancer diagnosis. Further details of the regression process are given in the Methods.
b) Pie-chart illustrating the heavy class-imbalance in the dataset.
@@
c) Class-balanced accuracy vs. training set length, comparing photonic network against software benchmarks. Test set is fixed at 1015 images. Class-balanced accuracy penalises incorrect classification on rarer classes more heavily, ensuring performance is not dominated by majority classes. Photonic network scores 53.7% max accuracy.
@@
Network laser performance relative to software benchmarks is assessed in figures 3g,h,i). Classification accuracy vs. training set length is assessed for single- and multi-layer network laser schemes against a range of software neural networks. Five software approaches are considered for all tasks: logistic regression on the raw images (a direct control, effectively bypassing our physical system), a multi-layer perceptron (300100 neurons), a LeNet5 CNN (22 filters), a larger CNN (96 filters), and the much more powerful pre-trained ‘EfficientNetV2-B0’[51, 52] CNN (3040 filters, 7.9 million parameters). For the harder BreaKHis 400X task with limited data we also benchmark using a far larger and more powerful pre-trained vision transformer network[39, 56, 7] (‘ViT_b_16’) with 86 million parameters, representing a large modern software vision network. Both ‘EfficientNetV2-B0’ and the vision transformer were trained using a transfer learning approach - initialised with pre-trained ImageNet weights, which are then updated during training. Full details of the software networks are provided in the Methods. We also benchmark against a ‘quasi-isomorphic’ software network: an initial convolutional filter layer of 10 44 filters followed by 30,000 randomly weighted ReLu neurons and a final logistic regression layer, mirroring the structure of our system. Results are shown in supplementary information.
Few-shot/low-data performance The photonic network laser excels when training data is scarce, a regime where conventional AI can struggle. The network outperforms all considered software models in the MNIST-digits and Fashion-MNIST tasks on training sets smaller than 5,000 images (digits), 15,000 images (fashion). On the harder biomedical BreaKHis task the network outperforms the same software benchmarks, including the powerful EfficientNet CNN, across all training lengths. Only the state-of-the-art vision transformer outperforms the photonic network at longer training lengths, with the photonic network outperforming even the vision transformer up to 200 training images - a substantial result.
We score 77.7% accuracy on BreaKHis from just 10 training images (5 examples per class), far higher than the software models. This highlights the effective learning provided by the nonlinear photonic dynamics and mode competition of the network laser. Here we compare performance when using only the raw training set. Software approaches can achieve higher scores, including on the BreaKHis task[57], but this requires costly extra steps including artificial augmentation of the training set, and extensive fine-tuning of model weights, which we do not perform in our photonic scheme.
@@
The neuromorphic capabilities of the photonic network are not restricted to feature detection and classification. Here, we demonstrate the ability of the network to spatially map regions of interest in biomedical images - drawing masks over skin lesions in colour photographic images from the HAM10000/ISIC 2018 dataset[58, 55] (fig. 4a, 10015 images of seven different classes of skin lesion). This task, termed segmentation, demonstrates that while the hyperspectral output of our network only contains information on the wavelength and one-dimensional spatial position of the lasing response, it is capable of physically encoding and processing complex two-dimensional spatial information - required for segmentation and other spatial processing. The task here is challenging: lesions are non-uniform in shape, a variety of colours and intensities, and many images contain darker non-lesion regions such as hairs, moles, and areas of uneven lighting. We combine two distinct tasks on the HAM10k dataset: class-imbalanced diagnostic classification, separating lesions into seven classes, and segmentation, drawing spatial masks to indicate lesion location. This two-fold processing has potential future impact in automated/semi-automated combined diagnosis and surgery systems. We only need to input the HAM10k images to our photonic network once to achieve both tasks, simply retraining our computationally-cheap linear output regression layer to change between diagnosis and segmentation.
@@
Class-imbalanced diagnosis Here we diagnose lesion images into seven classes. This is made highly challenging by the heavy class-imbalance of the dataset - illustrated in figure 4b). Machine learning typically struggles with class imbalance, tending to incorrectly classify rare classes as examples of the more common classes. Overcoming this requires costly techniques such as data augmentation, e.g. generating many distorted copies of training examples and using them to artificially extend the training set. The HAM10000 task is heavily imbalanced, with the ‘NV’ class representing 66.9% of the dataset, and four of the seven classes each representing 5.1% or less. Figure 4c) shows class-balanced classification accuracy against training set length (class-balanced accuracy penalises incorrect guesses on rare classes more heavily), with the photonic network scoring 53.7% max accuracy, outperforming all software benchmarks other than the vision transformer across all training lengths, and outperforming the vision transformer at training lengths up to 500 images. This demonstrates that our photonic system exhibits strong advantages on both few-shot learning and heavily-imbalanced datasets - a hard problem with real-world benefits.
@@
Biomedical segmentation For segmentation, we use the same hyperspectral responses from the lesion images which we use for classification. Here, rather than using the diagnostic class as training/testing labels, and logistic regression as an output layer - we use the ground-truth segmentation maps as training/testing labels and ridge regression as an output layer (as the segmentation task has continuous output, rather than discrete labels). Otherwise, the procedure follows that used for classification. Figure 4d) compares three segmentation metric scores across the test set: DICE (F1 score for segmentation), Jaccard (intersection over union) and pixel accuracy (correct pixels over incorrect). DICE and Jaccard provide best indication of segmentation quality. Pixel accuracy scores tend to be distorted by whether the correct ground-truth mask occupies a small or large fraction of the total image. The photonic network scores highest on all metrics, highlighting the ability of the network to implement spatial processing.
@@ Conclusions
This work highlights the benefit of introducing complex heterogeneous physical nonlinearities to photonic computing systems, particularly the combination of inhibitory and excitatory dynamics and spatially-distributed disordered mode profiles. By leveraging the ability of the disordered lasing network to provide heterogeneous nonlinear emission in both spatial and spectral domains alongside parallel feature extraction, we perform a range of vision tasks including feature detection, image classification, and segmentation, with promising performance relative to software approaches when learning with scarce training data. This bears interesting links with prior work exploring heterogeneous nonlinearity in spiking neural networks [23] and nanomagnetic systems [59], suggesting that engineering heterogeneous nonlinear dynamics across a broad range of photonic and non-photonic systems may provide fertile ground for the community to explore – with potential impact on biomedical sensing, edge computing, and beyond. As the lasing modes responsible for feature detection spatially overlap, we provide parallel feature detection of multiple image features in a single network with a small on-chip footprint (with additional off-chip optical components). In contrast, existing physical convolutional hardware schemes typically detect one feature per physical repeat of the device, hence the spatial footprint, fabrication complexity and energy consumption typically scale with the number of features detected.
Strong learning in the few-shot/low-data regime with scarce training examples is a hard and open challenge in machine learning. The performance of our photonic network in this regime, including on hard biomedical tasks such as BreaKHis and HAM10k, provides valuable evidence that physics-based neuromorphic systems can handle complex, data-constrained learning tasks. These results bring physical neuromorphic schemes closer to offering viable alternatives to conventional software-based approaches. Our good performance here extends to heavily class-imbalanced datasets, which are similarly challenging for conventional software AI. One of the key proposed applications for neuromorphic computing hardware is ‘edge computing’, where remotely deployed hardware performs computation directly at the point of sensing. While edge computing is often (though increasingly not solely[40, 41]) limited to inference, our system shows promise for edge deployments where both training and inference are performed, which requires adapting on-the-fly to incoming data and changing environmental conditions without access to large offline/cloud-based training data. Here, it is highly advantageous to perform well with short and imbalanced training sets, as locally collected data will typically be far from the ideal large and class-balanced training sets where software AI is most comfortable.
@@
RS, KM, HS, JD, AF, DS, and WKN designed and developed the network laser concept.
@@
JCG, WKN, JD, and AF wrote the manuscript with contributions from all authors.
@@ Acknowledgements
This work was supported by a Royal Academy of Engineering Research Fellowship awarded to JCG.
@@
JCG, RS and JD were supported by the Imperial College London President’s Excellence Fund for Frontier Research.
JCG was supported by the ERC Starting Grant MORPHON.
@@
TVR, DS, MB and RS acknowledge support from the Engineering and Physical Sciences Research Council (EPSRC), grant number EP/T027258/1.
WKN acknowledges research support through a President’s PhD Scholarship from Imperial College London.
This work was supported by funding from the EPSRC Impact Acceleration Account awarded to JCG and RS.
This work was supported by the Imperial College London DT-Prime scheme awarded to JCG, RS and MB.
@@
TF was supported by an EPSRC doctoral training programme PhD award.
@@
We thank David Mack and Stefano Vezzoli for excellent laboratory management, and Stephen Cussell and David Bowler for excellent technical workshop services at Imperial College London.
@@
Test image of ‘Mandrill’ was taken from the Test Image Collection at https://www.hlevkin.com/hlevkin/06testimages.htm, also available in the USC-SIPI image database at:
https://sipi.usc.edu/database/database.php.
JCG, WKN, JD, TVR, DS, and KM declare their commitment to responsible scientific conduct and explicitly oppose any dual-use or otherwise harmful application of the findings presented in this work. Specifically, we strictly oppose any activities pertaining to defence, arms or military applications.
@@
The authors declare the following competing interests: WKN, JD, AF, TVR, DS, KS, WB, MB, KM, RS, JCG are founders or shareholders of Rayd Technologies, a spinout company formed to commercialise the technology described in this work. A patent covering related subject matter has been published (WO2025114710A1).
Data and code availability
The data used to generate the figures in this manuscript, code implementing logistic regression classification, and a set of hyperspectra for the BreaKHis 400X task are available at https://doi.org/10.6084/m9.figshare.31209973. Further hyperspectral data is available from the corresponding author on reasonable request. The experiment automation frameworks used are open-source: https://puzzlepiece.readthedocs.io/, https://pzp-hardware.readthedocs.io/
@@ Methods
The network was designed by using a randomly-distributed point grid pattern on a plane and generating a Voronoi diagram[29]. A Voronoi diagram is generated by forming areas around stochastically placed points that contain all points closer to a certain point than any other point. The edges of these areas form the network design. Each vertex of the designed network is formed by three waveguides. The network is cropped to a circular area with 150 m diameter, and the density of the point patterns is varied to generate waveguide graph-edges with 350 nm width and an average length of 5 m.
@@
The edge-detection and digit classification experiments are performed in an optical microscopy setup with a spatially-structured pump (via DMD) to excite the network laser structure. A schematic of the setup is shown in the supplementary information, fig. S1). The pump laser is a 633 nm femtosecond pulsed laser (Light Conversion PHAROS with ORPHEUS-VIS Optical Parametric Amplifier (OPA) and LYRA harmonic generator, 200 fs pulse duration, 1 kHz repetition rate) is first expanded through a telescope and shaped by reflecting on a digital micromirror device (DMD, ViALUX V-650L). The DMD consists of programmable mirror pixels which shape the incident uniform beam into any pixelated black-and-white image (i.e. training images). For the edge detection and MNIST classification schemes, we use the DMD only as a binary modulator, with the DMD controller thresholding the provided images at the mid-way point in brightness. For the projection of Fashion-MNIST and BreaKHis 400X images we employ a quasi-greyscale scheme, using ordered dithering with a blue noise pattern, with multiple binary DMD pixels corresponding to one greyscale data pixel. The power of the shaped pump patterns is modulated by a motorised continuously variable neutral density (ND) filter, and the exact power received on the sample is monitored by a power meter (Thorlabs, S120VC Si photodiode, 200 - 1100 nm) on the split path reflected from a glass slide inserted. The spatially-structured pump is focused on the sample through a objective (Olympus Plan N 20x with a 0.4 numerical aperture (NA)). The lasing emission from the network is collected through the same objective, filtered (Thorlabs FELH0750, long-pass filter), and spatially and spectrally analysed using a grating spectrometer (Princeton Instruments SpectraPro HRS-300) equipped with a g/mm visible grating without an entrance slit. The signals are measured by a 2D charge-coupled device camera (CCD, Princeton Instruments Pixis 256) with a spatial resolution of m.
@@
The feature maps of the “Cameraman”, “CL”, MNIST, Fashion-MNIST, and BreakHis images (and “street” image in the supplementary) are generated from raster-scanning the binarised images across the kernels with a size of 4 4, and a stride of 1. For the ease of the data collection process and minimising redundant data collections, a set of hyperspectral data from all possible 4 4 image slices (a total of 65,536 excitation patterns) is collected sequentially. The feature maps of these images are then generated by picking the corresponding hyperspectra based on the slice windows. The same process applies when we generate the edge maps in the two-layer classification scheme.
The feature maps of the mandrill image (Fig. 2k) with the kernel sizes of 4 4, 7 7, and 11 11 are measured separately for a consistent comparison. The feature maps with RGB channels are generated by raster-scanning every image slice from the 256 246 pixel image (with the size of the kernel and stride of 2) on the network laser, comprising a total of 45,000 sequential acquisitions for each of the kernels.
@@
The lasing modes and intensities of the network are modelled using netSALT,[34] which extends steady-state ab initio laser theory (SALT)[48] for quantum graphs. The netSALT model was previously developed by several of the co-authors and a detailed derivation is provided in the Supplementary Information of Saxena et al.[34]. SALT solves the following partial differential equation, derived from the Maxwell Bloch equations that describe lasing two-level atoms interacting with light, to find the lasing mode profiles and their frequencies ,
where is the permittivity of the medium, is the central frequency of the lorentzian gain profile, is the gain width, is the lorentzian gain at the mode frequency , is the spatially varying pump intensity, and is the population inversion that is reduced from due to mode competition. NetSALT computes the lasing from networks by solving the one-dimensional SALT equation on the edges of the network and applying continuity of fields and flux at the nodes of the network[34].
The designed network is represented as a graph, with the edges subdivided to overlap with an 88 grid to correspond to the kernel. The edges of the network have a complex permittivity when unpumped, a lorentzian gain profile centered at normalised frequency with a width , and are pumped up to a normalised power (For a detailed physical explanation of the parameters, see Ref. 34). A separate simulation is performed for each 88 binary kernel, by pumping only the edges falling under the illuminated pixels. The passive (without pumping) modes of the network are first identified in the complex frequency plane , and these are the same for all kernels. For each kernel, the modes are tracked in the complex plane as the pump power is increased, allowing to compare between the same modes across different kernels. As a mode reaches the real frequency plane [Im], it attains threshold and starts lasing. Beyond this pump power, the frequency and shape of the mode do not change but mode competition between the lasing modes modifies their lasing intensities. They further deplete the gain available to the remaining modes, increasing their thresholds. The mode frequencies, thresholds and intensities found by netSALT incorporate the mode competition, gain depletion and nonlinearity.
@@
Within the training sets, we set aside a fixed 10% validation split to monitor generalization and detect overfitting. For CNN models, weights are initialized using Xavier/Glorot initialization (biases set to zero). We train using the Adam optimizer with learning rate , batch size 128, for 20 epochs, using cross-entropy loss. To ensure reproducibility, we report the full list of trial seeds and use a fixed dataset subsampling/splitting seed; results are aggregated across 10 independent trials.
All software CNN baseline training was performed on an NVIDIA GeForce RTX 4070 GPU. To address overfitting, validation accuracy vs test accuracy as a function of epoch (mean standard deviation across trials) for each dataset/model pair (MNIST-LeNet-5, MNIST-96-filter CNN, FashionMNIST-LeNet-5, FashionMNIST-96-filter CNN) are monitored. The plots of validation/test accuracy over training epochs are shown in SI Fig. S25, and shows that validation and test accuracies track well during training, reaching a plateau without diverging. Overfitting issues are not observed.
CNN training parameters
@@
| Parameter | Value |
| Total samples () | 40,000 |
| Training samples () | 36,000 |
| Test samples () | 4,000 |
| Validation fraction () | 0.10 |
| Epochs | 20 |
| Batch size | 128 |
| Learning rate | |
| Optimizer | Adam |
| Number of trials () | 10 |
| Split seed | 12345 |
| Trial seeds | |
| Initialisation | Xavier (PyTorch) |
| Training hardware | NVIDIA 4070 GPU |